OpenStudy (anonymous):

Please help with this integral : e^(-x)/x^3 as x goes to infinity ... now i know the answer is infinity but i don't know how to solve it to get to the answer =\

hartnn (hartnn):

integral or limit??

OpenStudy (anonymous):

integral =)

hartnn (hartnn):

integral from 0 to infinity or -infinity to infinity?

OpenStudy (anonymous):

0 to infinity

hartnn (hartnn):

have u tried uv(product) rule by making the function e^(-x)x^(-3)?

OpenStudy (anonymous):

\[\int_{0}^{\infty} \frac{e^{-x}}{x^3} \text{d}x\]?

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

i think integration by parts will work here...as hartnn said are u familiar with integration by parts?

OpenStudy (anonymous):

yes i am .. but i got integral of e^(-x)*lnx eventually which i couldn't solve =\

OpenStudy (anonymous):

howahhh...i found something finally...with integration by parts u will get : \[\int_{0}^{\infty} \frac{e^{-x}}{x^3} \text{d}x=-\frac{e^{-x}}{2x^2} \mid_0^\infty+\frac{e^{-x}}{2x} \mid_0^\infty+\int_{0}^{\infty} \frac{e^{-x}}{x} \text{d}x\]check this then i will continue :)

OpenStudy (anonymous):

lol... @PeX are u there?

OpenStudy (anonymous):

there is a trick to do that

OpenStudy (anonymous):

i think u should continue by taking 1/x term as the derivative unitill it reaches the power of 1/x3 then bring this integral to the left side and add with e^(-x)/x^3

OpenStudy (anonymous):

nope it cancels everything...and we will have the same thing

OpenStudy (anonymous):

\[\int_{0}^{\infty} \frac{e^{-x}}{x^3} \text{d}x=-\frac{e^{-x}}{2x^2} \mid_0^\infty+\frac{e^{-x}}{2x} \mid_0^\infty+\int_{0}^{\infty} \frac{e^{-x}}{x} \text{d}x\\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ =\lim_{x \rightarrow 0} (\frac{1}{2x^2}-\frac{1}{2x})+\int_{0}^{\infty} \frac{e^{-x}}{x} \text{d}x=+\infty+\text{Something >0}=+\infty \]

OpenStudy (anonymous):

\[\int\limits_{}^{}e^-x/x=?\] now let 1/x=lnt e^1/x=t we have 1/x=lnt and e^-x=t substitute \[\int\limits_{}^{}tlnt\] = \[(t^2/2)lnt -\int\limits_{}^{}t/2\] \[t^2lnt/2-t^/4\]

OpenStudy (anonymous):

there should be -t^2/4 last expression

OpenStudy (anonymous):

sorry i tried it's extremely wrong

OpenStudy (anonymous):

Thank you very much !!! you are the math messiah =)

OpenStudy (anonymous):

yw...:)